# Attention (as Discrete-Time Markov) Chains

### v1

@@ 5.2 Linear Probing of TokenRank
5.3 Improving Self-Attention Guidance (SAG)
To evaluate TokenRank on a downstream task, we integrate it in SAG (Hong et al., 2023), an approach that improves the fidelity and the quality of unconditional image generation with denoising diffusion models. SAG uses the self-attention weight matrices to form a mask that indicates important spatial tokens. It then adversarially blurs the masked regions and drives the denoising process away from it, creating better images as a result. To find out where important features exist, Hong et al. (2023) average over the head dimension for a specific attention layer, followed by a summing each of the columns (Section 4.1). We compare the quality and diversity of generated images when using the TokenRank instead of their proposed Column-sum strategy. We compute the IS (Salimans et al., 2016), FID (Heusel et al., 2017), and KID (Bińkowski et al., 2018) metrics over 50K generated images. The selected metrics offer complementary information in terms of quality and diversity and resemblance to ground-truth distribution. We refer to the supplementary for further details. Results are reported in Table 3 and Figure 12. Using TokenRank improves the generated images significantly in both quality and diversity compared to the original SAG, showcasing TokenRank’s ability to rank globally important tokens in the sequence, and improve a downstream generative task.
@@ 5.4 Masking Most Influential Tokens
Table 4: Results for masking most influential tokens. We present the area under curve (AUC) metric for the accuracy normalized by the original model accuracy and average over various models of a selected model family with model-specific results presented in the supplementary. Using TokenRank to mask out important tokens consistently results in larger accuracy drops than other strategies.
@@
Inspired by standard faithfulness experiments (Mehri et al., 2024), where changes in the performance of a downstream classifier are observed through progressively occluding input features, we designed a new experiment that targets evaluating strategies that measure the global importance of tokens per individual attention head. Specifically, the most influential tokens in the sequence are masked out progressively by zeroing the corresponding columns in the attention matrices and the resulting classification accuracy drop is measured. We evaluate the accuracy degradation of several vision transformer-based classifiers over 5000 randomly selected images from ImageNet with a fixed seed. To determine the order of tokens to be masked, we use the TokenRank vector and compare it to the following baselines: randomly selecting a token (“Rand. Token”), using the column of the token corresponding to the center patch of the image (“Center Token”), the per column-sum (“Column Sum”), and using the CLS Token. Note that we do not mask global tokens, i.e., the CLS token and the registers for DINOv2, as this results in a massive drop in performance rendering the comparisons less useful. We refer to the supplementary for further details.
Results are reported in Table 4, illustrating that classification accuracy degrades faster using TokenRank than other baselines for several pre-trained transformers. This validates the hypothesis that TokenRank captures important global information more faithfully since it also models higher-order effects. Another interesting observation is that transformers with a more structured features (e.g. DINOv2 with registers (Darcet et al., 2024)) tend to benefit more from TokenRank, which suggests it can also be potentially used for quantifying this aspect of a trained transformer. We hypothesize that this is due to TokenRank’s tendency to sharpen the input signal (Figure 4), potentially reinforcing the effect of noise or artifacts for an unstructured feature space.
@@ 6 Conclusion
In this paper, we proposed a novel interpretation of the attention map as a DTMC. By treating such maps as stochastic matrices, we showed that taking into account higher-order interactions between tokens results in better performance on various downstream tasks. We further showed the stationary vector TokenRank can help depict global incoming and outgoing attention within a single head. We envision that this framework opens up new possibilities for analyzing, understanding, and modeling the attention operation, and may be used in conjunction with other end-to-end explainability tools to enhance understanding of visual transformers. This can also serve various other domains such as video, audio, motion, as well as natural language processing.
@@
Appendix A Appendix
A.1 Implementation Details
A.1.1 ImageNet Zero-Shot Segmentation
To establish fair comparisons, we followed the exact procedure described by Helbling et al. (2025) using the code from https://github.com/helblazer811/ConceptAttention. We simplify all categories to a single word, allowing methods conditioned on text to select the corresponding token. We use the timestep-distilled Flux-Schnell DiT (Labs, 2024) and add Gaussian noise to each of the images to the normalized timestep $t=0.5$, resulting in two backward steps for denoising the image. We use the last 10 layers of the multi-stream attention blocks. The attention matrices from all aforementioned layers and timesteps are averaged. For the heads, we use the $\lambda_{2}$ weighting for averaging. For multi-bounce attention, we use the column-select formulation, i.e. we replace $A$ with $A_{l}$ and transpose as described in Section 4.2 of the paper.
A.1.2 Linear Probing of TokenRank
We generate attention maps for all layers, heads, and images for the ImageNet (Deng et al., 2009) subset Imagenette (Howard, 2019) for DINOv1, DINOv2, and CLIP with a VIT/B model. Then, we compute attention visualizations with column sum, TokenRank, row-selecting the center token, and row-selecting the CLS token. Finally, we train a linear classifier with the commonly used train and test split and with the attention visualizations of all layers and heads as input, optimized with the SGD optimizer with a set of learning rates ({$51e-$ 5, $52e-$ 5, $55e-$ 5, $41e-$ 4, $42e-$ 4, $45e-$ 4, $31e-$ 3})) for 20 epochs, and we choose the best performing model, similar to Oquab et al. (2024). We resize and center crop the images to get $905\times 905$ ($901\times 901$) attention maps for DINOv2 with registers (other models). We exclude the CLS token column for classification.
A.1.3 Improving Self Attention Guidance (SAG)
For the SAG experiments, we build on top of the code of (Hong et al., 2023) provided in https://github.com/SusungHong/Self-Attention-Guidance. We use SD1.5 (Rombach et al., 2022) as the foundational model and use the 2nd last decoder block’s last self-attention layer for masking. We empirically observed it consistently resulted in the best results for both SAG and SAG+TokenRank, as opposed to the usage of the bottleneck layer in the comparison experiments in the original SAG, which is in line with the observation by Hong et al. (2023).
@@ A.1.3 Improving Self Attention Guidance (SAG)
A.1.4 Masking Most Influential Tokens
@@
For all transformers, we do not allow masking the CLS token in our evaluations, as it consistently ranks as the top most important token for almost all heads and layers, and discarding it yields an extreme drop in accuracy, which does not allow for comparisons of the alternative approaches. For the same reason, we also only use the first half of the available attention layers for each model. This allows for not disturbing the output space too much and as a result accuracy performance degrades gracefully. Refer to Section A.6.2 for further justification of this choice. Moreover, we mask out the same number of tokens for all heads simultaneously. For DINO family type, results in the main paper are averaged over the ViT-B/16, ViT-B/32, ViT-S/16, and ViT-S/32 architectures. For DINOv2 family type, we use the w/ registers variation, and average over the ViT-B/14, ViT-S/14 and ViT-L/14 architectures. For both DINO and DINOv2 family types, we use pretrained linear classifier heads provided with the original publications. We average ViT/B and ViT/L for the transformer models trained in a supervised way. For CLIP family type, we average over the ViT-B/32, ViT-L/14 and ViT-L/14/336 architectures. We use the template prompt “A photo of a <class>" to compute per-class text features. Then, we compute a dot-product with the image features and these per-class features to determine the classification result. Masking is performed by setting the entire column corresponding to a token to $\infty$- ∞ prior to softmax. For a more in-depth view of individual model accuracy degradation, see Section A.6.
A.2 Ablation and Hyperparameters for Zero-Shot Segmentation
@@
See Table 5 for zero-shot ImageNet segmentation results using the same experimental setup as in the main paper, with different hyper parameter choices. The bottom of Table 5 shows “outgoing ($n=2$)*”, corresponding to the result shown in the main paper, leveraging the 2nd bounce from a one-hot column vector and utilizing $\lambda_{2}$ weighting. “row-select ($n=1$)” and “column-select ($n=1$)” are the naive first bounce operation also shown in the main paper. “single-channel” uses the single channel attention blocks in FLUX rather than the dual-channel blocks. “incoming ($n=2$)” uses the incoming attention formulation of multi-bounce-attention (i.e. 2nd bounce from a one-hot row vector). “incoming ($n=3$)” is the same but with 3 bounces. “outgoing ($n=3$)” uses a one-hot column vector with 3 bounces. “w/o $\lambda_{2}$” does not use the weighted $\lambda_{2}$ scheme. Notice the performance for outgoing attention degrades rapidly with more bounces (“outgoing ($n=2$)*”, “outgoing ($n=3$)”) compared to incoming attention (“incoming ($n=2$)”, “incoming ($n=3$)”). This is because for outgoing attention, the steady-state eventually converges into emphasizing important “hubs”, which usually tends to be background information. However, in early bounces ($n=1,2$) it out-performs the incoming attention, and we speculate this is because it captures better the target of this task: we are interested in finding out which image tokens “attend-to” the text token of interest (i.e. a many-to-one relationship).
@@
w/o $\lambda_{2}$ 84.00 70.02 94.28
outgoing ($n=2$)* 84.12 70.20 94.29
A.3 Used Compute
A.4 More Segmentation Results
See Figures 13 and 14 for more zero-shot segmentation results on ImageNet with comparisons to row- and column select and state-of-the-art results using ConceptAttention (Helbling et al., 2025). We argue that our raw segmentation masks are cleaner and more precise, which can be qualitatively measured by the threshold-agnostic mAP metric presented in the main paper. Additionally, in Figure 6 we show our multi-bounce attention can be used for segmenting different parts of an image by setting the initial one-hot vector to the corresponding text token. The text prompt used to create the image (FLUX-Schnell) was “cute black cat standing up wearing red boots and a bow tie, photorealistic, masterpiece, macro wildlife photography, dewy grass, closeup low angle, wildflowers, sun, shadows, depth of field, desaturated, film grain, low quality”.

### v2

@@ 5.3 Improving Self-Attention Guidance (SAG)
5.4 Improving DiffSeg with TokenRank
Table 3: COCO-Stuff-27 benchmark. Using TokenRank to sample anchors for initial proposals in DiffSeg (Tian et al., 2024) results in significant improvements over the benchmark.
Method mACC $\uparrow$ mIoU $\uparrow$
Uniform Grid 72.50 43.60
TokenRank Grid 84.97 44.87
We incorporate TokenRank into DiffSeg (Tian et al., 2024), an existing segmentation approach that merges self-attention matrices using KL-divergence to measure similarity between upsampled attention maps. Specifically, in the original study the authors used a uniform grid to initialize the anchors used as seeds for their proposal algorithm. Instead, we sample anchors according to their TokenRank importance, with suppression to avoid repeatedly selecting the same location. This better grid strategy leads to substantial improvements over the original approach using the COCO-Stuff-27 benchmark as can be seen in Table˜3. This experiment further shows our framework is orthogonal to other solutions leveraging self-attention for solving downstream tasks.
@@ 5.5 Masking Most Influential Tokens
Inspired by standard faithfulness experiments (Mehri et al., 2024), where changes in the performance of a downstream classifier are observed through progressively occluding input features, we designed a new experiment that targets evaluating strategies that measure the global importance of tokens per individual attention head. Specifically, the most influential tokens in the sequence are masked out progressively by zeroing the corresponding columns in the attention matrices and the resulting classification accuracy drop is measured. We evaluate the accuracy degradation of several vision transformer-based classifiers over 5000 randomly selected images from ImageNet with a fixed seed. To determine the order of tokens to be masked, we use the TokenRank vector and compare it to the following baselines: randomly selecting a token (“Rand. Token”), using the column of the token corresponding to the center patch of the image (“Center Token”), the per column-sum (“Column Sum”), and using the CLS Token. Note that we do not mask global tokens, i.e., the CLS token and the registers for DINOv2, as this results in a massive drop in performance rendering the comparisons less useful. We refer to Appendix˜A for further details.
Table 4: Results for masking most influential tokens. We present the area under curve (AUC) metric for the accuracy normalized by the original model accuracy and average over various models of a selected model family with model-specific results presented in Appendix˜D. Using TokenRank to mask out important tokens consistently results in larger accuracy drops than other strategies.
@@
Results are reported in Table˜4, illustrating that classification accuracy degrades faster using TokenRank than other baselines for several pre-trained transformers. This validates the hypothesis that TokenRank captures important global information more faithfully since it also models higher-order effects. Another interesting observation is that transformers with a more structured features (e.g. DINOv2 with registers (Darcet et al., 2024)) tend to benefit more from TokenRank, which suggests it can also be potentially used for quantifying this aspect of a trained transformer. We hypothesize that this is due to TokenRank’s tendency to sharpen the input signal (Figure˜4), potentially reinforcing the effect of noise or artifacts for an unstructured feature space.
@@ 6 Conclusion
In this paper, we proposed a novel interpretation of the attention map as a DTMC. By treating such maps as stochastic matrices, we showed that taking into account higher-order interactions between tokens results in better performance on various downstream tasks. We further showed the stationary vector TokenRank can help depict global incoming and outgoing attention within a single head.
Limitations and future work.
Our approach is limited to square matrices (e.g. self-attention / hybrid-attention blocks). Therefore, cross-attention blocks do not naturally fit into our Markov chain formulation due to the existence of non-accessible states. There could potentially be a way to resolve this by introducing dummy states with uniformly distributed transition probability. Additionally, computing the second eigenvalues $\lambda_{2}$ is computationally heavy. Computing multiple bounces or the TokenRank on the other hand is quite performant, and converges typically after 10 to 20 iterations. Finally, we observe that the performance gains with TokenRank are larger for transformers where the attention map is more structured. We envision that this framework opens up new possibilities for analyzing, understanding, and modeling the attention operation, and may be used in conjunction with other end-to-end explainability tools to enhance understanding of visual transformers. This can also serve various other domains such as video, audio, motion, as well as natural language processing.
@@
Supplementary Material
Appendix A Implementation Details
A.1 ImageNet Zero-Shot Segmentation
To establish fair comparisons, we followed the exact procedure described by Helbling et al. (2025) using the code from https://github.com/helblazer811/ConceptAttention. We simplify all categories to a single word, allowing methods conditioned on text to select the corresponding token. We use the timestep-distilled Flux-Schnell DiT (Labs, 2024) and add Gaussian noise to each of the images to the normalized timestep $t=0.5$, resulting in two backward steps for denoising the image. We use the last 10 layers of the multi-stream attention blocks. The attention matrices from all aforementioned layers and timesteps are averaged. For the heads, we use the $\lambda_{2}$ weighting for averaging. For multi-bounce attention, we use the column-select formulation, i.e. we normalize the columns and transpose $A$ as described in Section˜4.2 of the paper.
A.2 Improving Self Attention Guidance (SAG)
For the SAG experiments, we build on top of the code of (Hong et al., 2023) provided in https://github.com/SusungHong/Self-Attention-Guidance. We use SD1.5 (Rombach et al., 2022) as the foundational model and use the second last decoder block’s last self-attention layer for masking. We empirically observed it consistently resulted in the best results for both SAG and SAG+TokenRank, as opposed to the usage of the bottleneck layer in the comparison experiments in the original SAG, which is in line with the observation by Hong et al. (2023).
@@ A.2 Improving Self Attention Guidance (SAG)
A.3 Improving DiffSeg with TokenRank
We used the official code of Tian et al. (2024) provided in https://github.com/google/diffseg. DiffSeg aggregates attention weights from all layers, where deeper layers are up-sampled. The aggregation is also a Markov Chain (see Section˜4.1) and, therefore, TokenRank can be computed once per image. To create the TokenRank grid, we simply compute TokenRank and sort it by descending order, and greedily select the highest ranked token that is at least 5 units away from all currently selected tokens to act as suppression which avoid oversampling the same region. If not enough anchors can be selected in this way, we remove the suppression and select the rest of the tokens by only considering their TokenRank. We use the same total amount of anchors as the baseline for fairness.
A.4 Masking Most Influential Tokens
@@
For all transformers, we do not allow masking the CLS token in our evaluations, as it consistently ranks as the top most important token for almost all heads and layers, and discarding it yields an extreme drop in accuracy, which does not allow for comparisons of the alternative approaches. For the same reason, we also only use the first half of the available attention layers for each model. This allows for not disturbing the output space too much and as a result accuracy performance degrades gracefully. Refer to Section˜D.2 for further justification of this choice. Moreover, we mask out the same number of tokens for all heads simultaneously. For DINO family type, results in the main paper are averaged over the ViT-B/16, ViT-B/32, ViT-S/16, and ViT-S/32 architectures. For DINOv2 family type, we use the w/ registers variation, and average over the ViT-B/14, ViT-S/14 and ViT-L/14 architectures. For both DINO and DINOv2 family types, we use pretrained linear classifier heads provided with the original publications. We average ViT/B and ViT/L for the transformer models trained in a supervised way. For CLIP family type, we average over the ViT-B/32, ViT-L/14 and ViT-L/14/336 architectures. We use the template prompt “A photo of a <class>" to compute per-class text features. Then, we compute a dot-product with the image features and these per-class features to determine the classification result. Masking is performed by setting the entire column corresponding to a token to -$\infty$ prior to softmax. For a more in-depth view of individual model accuracy degradation, see Appendix˜D.
Appendix B Linear Probing of TokenRank
Accuracy DINOv1 DINOv2 CLIP
Rand. Token $67.33\pm 3.12$ $63.32\pm 6.04$ $57.65\pm 2.66$
Center Token $66.32\pm 3.08$ $75.34\pm 1.85$ $68.42\pm 2.66$
Column Sum $75.64\pm 2.74$ $90.76\pm 2.21$ $73.73\pm 2.98$
TokenRank $\mathbf{77.31\pm 2.50}$ $\mathbf{92.73\pm 1.51}$ $\mathbf{73.88\pm 2.86}$
CLS Token $81.53\pm 2.44$ $94.07\pm 0.97$ $72.46\pm 3.40$
To quantitatively evaluate that TokenRank produces sharper maps and captures the global information contained in the attention matrices, we train a linear classifier on top of all proposed attention visualizations, inspired by a previous similar analysis (Liu et al., 2024). For this purpose, we extract attention visualizations for all layers, heads, and images for the ImageNet (Deng et al., 2009) subset Imagenette (Howard, 2019) for three foundation models without considering the CLS token. The results in Table˜5 show that TokenRank visualizations result in a cleaner signal for image classification than previous approaches that only consider one bounce of attention. It is on par with row-selecting the CLS token, which is an upper bound because it was specifically trained to capture class concepts for the DINO model family. Note error bars are reported over 5 different runs each with a different seed. We performed a Wilcoxon test with the null hypothesis that the competing column sum operation results in the same accuracy as TokenRank and the one-sided alternative hypothesis that the accuracy is lower. We reject the null hypothesis for DINOv1 with p=0.03 and for DINOv2 with p=0.03.
We generate attention maps for all layers, heads, and images for the ImageNet (Deng et al., 2009) subset Imagenette (Howard, 2019) for DINOv1, DINOv2, and CLIP with a VIT/B model. Then, we compute attention visualizations with column sum, TokenRank, row-selecting the center token, and row-selecting the CLS token. Finally, we train a linear classifier with the commonly used train and test split and with the attention visualizations of all layers and heads as input, optimized with the SGD optimizer with a set of learning rates ({1e51e-5, 2e52e-5, 5e55e-5, 1e41e-4, 2e42e-4, 5e45e-4, 1e31e-3})) for 20 epochs, and we choose the best performing model, similar to Oquab et al. (2024). We resize and center crop the images to get $905\times 905$ ($901\times 901$) attention maps for DINOv2 with registers (other models). We exclude the CLS token column for classification.
Linear Probing of $\lambda_{2}$-Weighted Aggregated Heads
Additionally, we also train linear classifiers on top of the TokenRank vectors (incoming attention) per layer, where we use different weighting schemes for the head dimension. For this, we first compute the weighting coefficients per head and perform a weighted average of the heads. Then, we compute TokenRank for the aggregated results. We compare uniform weighting (simple average), random weighting, and $\lambda_{2}$ weighting. Results are presented in Table˜6: While random weighting is significantly worse, $\lambda_{2}$ weighting performs better than simple averaging, particularly for transformers with a well-structured latent space, such as DINOv2 with registers.
Method uniform random $\lambda_{2}$
CLIP $\mathbf{49.07}$ $44.07$ $\mathbf{49.07}$
DINOv1 $53.50$ $47.21$ $\mathbf{53.55}$
DINOv2 $71.62$ $50.29$ $\mathbf{72.36}$
Appendix C Additional Segmentation Results
C.1 Ablation and Hyperparameters
@@
See Table˜7 for zero-shot ImageNet segmentation results using the same experimental setup as in the main paper, with different hyper parameter choices. The bottom of Table˜7 shows “outgoing (n=2n=2)*”, corresponding to the result shown in the main paper, leveraging the second bounce from a one-hot column vector and utilizing $\lambda_{2}$ weighting. “row-select (n=1n=1)” and “column-select (n=1n=1)” are the naive first bounce operation also shown in the main paper. “single-channel” uses the single channel attention blocks in FLUX rather than the dual-channel blocks. “incoming (n=2n=2)” uses the incoming attention formulation of multi-bounce-attention (i.e. second bounce from a one-hot row vector). “incoming (n=3n=3)” is the same but with 3 bounces. “outgoing (n=3n=3)” uses a one-hot column vector with 3 bounces. “w/o $\lambda_{2}$” does not use the weighted $\lambda_{2}$ scheme. Notice the performance for outgoing attention degrades rapidly with more bounces (“outgoing (n=2n=2)*”, “outgoing (n=3n=3)”) compared to incoming attention (“incoming (n=2n=2)”, “incoming (n=3n=3)”). This is because for outgoing attention, the steady-state eventually converges into emphasizing important “hubs”, which usually tends to be background information. However, in early bounces ($n=1,2$) it out-performs the incoming attention, and we speculate this is because it captures better the target of this task: we are interested in finding out which image tokens “attend-to” the text token of interest (i.e. a many-to-one relationship).
We additionally report error bars in Table˜8 that were computed as standard deviations over 5 runs with different seeds.
@@
outgoing (n=2n=2)* 84.12 70.20 94.29
Table 8: ImageNet segmentation. Error bars for selected runs.
Method Acc mIoU mAP
Concept Attention $81.06\pm 0.01$ $66.02\pm 0.01$ $88.43\pm 0.00$
Ours no $\lambda_{2}$ $84.00\pm 0.02$ $69.99\pm 0.05$ $94.26\pm 0.04$
Ours $84.11\pm 0.03$ $70.19\pm 0.02$ $94.32\pm 0.03$
C.2 Further Qualitative Results
See Figures˜13 and 14 for more zero-shot segmentation results on ImageNet with comparisons to row- and column select and state-of-the-art results using ConceptAttention (Helbling et al., 2025). We argue that our raw segmentation masks are cleaner and more precise, which can be qualitatively measured by the threshold-agnostic mAP metric presented in the main paper. Additionally, in Figure˜6 we show our multi-bounce attention can be used for segmenting different parts of an image by setting the initial one-hot vector to the corresponding text token. The text prompt used to create the image (FLUX-Schnell) was “cute black cat standing up wearing red boots and a bow tie, photorealistic, masterpiece, macro wildlife photography, dewy grass, closeup low angle, wildflowers, sun, shadows, depth of field, desaturated, film grain, low quality”.
@@ C.2 Further Qualitative Results
Appendix D Additional Results for Masking Most Influential Tokens
D.1 Per-Model Results
In Figure˜7 we show the results of performing the most-influential token masking experiment described in the main paper for individual models. The area-under-curve (AUC) metric in the main paper is computed over the curves.
D.2 Per-Layer Results
In Figure˜8, we show the result of choosing progressively more number of layers to mask for the most-influential token masking experiment conducted in the main paper. “6 (Ours)” indicates what we used in the paper. Layers are selected starting from the furthest from the output. The results are shown for DINOv2 (ViT-S/14) w/ registers over the same dataset used in the main paper. However, similar trends were observed for all architectures. We opted for masking tokens from only the first 50% of the layers to avoid distorting the output space and a too steep performance drop (which happens for $n>6$), while still having an effect on the classification results for higher token mask percentage (which does not occur for $n<6$).
Appendix E Additional Results for $\lambda_{2}$ Weighting
E.1 Statistical Tests for $\lambda_{2}$ Weighting
In this section we provide more evidence that the magnitude of the second largest eigenvalue of the attention matrix correlates with the performance in both zero-shot segmentation (Section˜5.1) and linear probing (Appendix˜B), as presented in the main paper. We illustrated this effect via weighted averaging of heads. To test the significance of our observation, we perform a paired data statistical test (Wilcoxon test) for both tasks. For this purpose, we define the null hypothesis: Uniform and $\lambda_{2}$ weighting have the same performance. And the one-sided alternative hypothesis: Weighting results in a better performance. Results of can be seen in Table˜9. Rejecting the null hypothesis with $p<0.05$ supports our statements.
Table 9: Statistical test for $\lambda_{2}$ weighting. The results of the Wilcoxon test show that $\lambda_{2}$ weighting is significantly better than uniform weighting.
Semantic Segmentation Linear Probing
Metric Statistic p-value Model Statistic p-value
Acc 15.0 0.031 DINOv1 45.0 0.002
mIoU 15.0 0.031 DINOv2 42.0 0.010
mAP 15.0 0.031 CLIP 45.0 0.002
E.2 $\lambda_{2}$ Weighting for Masking Most Influential Tokens Experiment
We performed another token masking experiment similar to the one conducted in of the paper, only using TokenRank for determining the masking order. In this experiment, we compare using uniform weights per head (as in main paper), random head weights, and using $\lambda_{2}$ weighting. Results can be seen in Figure˜9. Perhaps unsurprisingly, randomly weighting the heads performs better than uniform weights, as the transformer is more sensitive to individual heads being fully or almost fully masked out, rather than gradually masking out all heads together. Importantly, results indicate that $\lambda_{2}$ weighting can be used for increasing the degradation in accuracy, suggesting it can serve as a useful tool for determining importance of heads. This was also observed to improve zero-shot segmentation results (see Appendix˜C).